if a is a negative number which of the following represents a postive number -3a, a\5,a-4, 2a-4
Answer by bobdabuildr(27)

if a is a negative number means -a
Plug -a into each of the answers to see which one it changes
-3(-a)=3a Multiplying two negatives makes a positive
-a/5=-a/5 Dividing one negative by a positive makes a negative
-a-4=-a-4 Subtracting a negative from negative makes a negative
2(-a)-4=-2a-4 Multiplying a negative and a positive makes a negative and subtracting a negative from a negative makes a negative